# 3cm3cm3cm3cm

### v1

@@ 3.3. The Bethe algebra for ${\mathfrak{gl}}_{d}$ and Frobenius algebras
for some $b_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$. According to [MR, Corollary 3.3] (see also [ChL2, Proposition 3.3]),
$b_{i}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$
for all $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$.
Now let
${\bf z}=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}.$
Define $\mu_{\bf z}\in{\mathfrak{gl}}_{p+m|q+n}^{*}$ by
@@
for $a_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ and $s\in{\mathbb{Z}}$.
Now consider the $(p+q+m+n)\times(p+q+m+n)$ matrix
${\mathcal{L}}_{{p+m|q+n}}^{\bf z}:=\Big{[}\Psi^{\bf z}\big{(}\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\big{)}\Big{]}_{i,j\in\mathbb{I}}=\Big{[}\delta_{i,j}(\partial_{z}-z_{i})-(-1)^{|i|}E^{i}_{j}(z)\Big{]}_{i,j\in\mathbb{I}}.$
Evidently, ${\mathcal{L}}_{{p+m|q+n}}^{\bf z}$ is an amply invertible Manin matrix of type ${\bf s}$ over the superalgebra $U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$ of pseudo-differential operators over $U({\mathfrak{gl}}_{p+m|q+n}[t])$. Just like Proposition 3.11, we have the following.
@@
${\rm Ber}\,^{{\bf s}^{\sigma}}\big{(}\big{(}{\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big{)}^{\sigma}\big{)}={\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right).$
@@
${\rm Ber}^{\bf s}\,\!\left({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\right)=\sum_{i=-\infty}^{p+m-q-n}b_{i}^{{\bf z}}(z)\partial_{z}^{i},$ (3.9)
where $b_{i}^{{\bf z}}(z):=\Psi^{\bf z}(b_{i})$. Let ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ be the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series $b_{i}^{{\bf z}}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. It is called the Bethe algebra of ${\mathfrak{gl}}_{p+m|q+n}$ with respect to ${\bf z}$. If we take $q=n=0$, we recover the Bethe algebra of ${\mathfrak{gl}}_{p+m}$ by virtue of Proposition 3.5. According to [MR, Corallary 3.7], the algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ is a commutative subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$. It also commutes with the Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ of ${\mathfrak{gl}}_{p+m|q+n}$ and hence acts on any weight space of a ${\mathfrak{gl}}_{p+m|q+n}$-module.
For any ${\mathcal{B}_{p+m|q+n}^{\bf z}}$-module $V$, let $\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}_{V}$ denote the image of the Bethe algebra ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ in ${\rm End}(V)$. We call $\big{(}{\mathcal{B}_{p+m|q+n}^{\bf z}}\big{)}_{V}$ the Bethe algebra of $V$ with respect to ${\bf z}$.
4. Duality for Bethe algebras
This section is devoted to establishing a duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ between the Bethe algebras $\mathcal{B}_{d}^{\bf w}\subseteq U({\mathfrak{gl}}_{d}[t])$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}\subseteq U({\mathfrak{gl}}_{p+m|q+n}[t])$ for any ${\bf w}\in{\mathbb{C}}^{d}$ and ${\bf z}\in{\mathbb{C}}^{p+q+m+n}$.
4.1. The actions of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$
Let $\mathcal{D}$ be the associative unital superalgebra generated by
$x^{a}_{i},\,\eta^{a}_{j},\,y^{a}_{r},\,\zeta^{a}_{s}$
as well as their derivatives
@@ 4. Duality for Bethe algebras
for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$ and $a=1,\ldots,d$. Here $x^{a}_{i}$, $y^{a}_{r}$, $\partial_{x^{a}_{i}}$ and $\partial_{y^{a}_{r}}$ are even while $\eta^{a}_{j}$, $\zeta^{a}_{s}$, $\partial_{\eta^{a}_{j}}$, and $\partial_{\zeta^{a}_{s}}$ are odd. The superalgebra $\mathcal{D}$ is called the Weyl superalgebra.
There is an anti-involution $\omega:\mathcal{D}\longrightarrow\mathcal{D}$ defined by
${\begin{array}[]{lll}x^{a}_{i}\mapsto\partial_{x^{a}_{i}},&\quad\partial_{x^{a}_{i}}\mapsto x^{a}_{i},\quad\eta^{a}_{j}\mapsto\partial_{\eta^{a}_{j}},&\quad\partial_{\eta^{a}_{j}}\mapsto\eta^{a}_{j},\\ y^{a}_{r}\mapsto\partial_{y^{a}_{r}},&\quad\partial_{y^{a}_{r}}\mapsto y^{a}_{r},\quad\zeta^{a}_{s}\mapsto\partial_{\zeta^{a}_{s}},&\quad\partial_{\zeta^{a}_{s}}\mapsto\zeta^{a}_{s},\end{array}}$ (4.1)
for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, and $1\leq s\leq q$. In light of Proposition 3.6, the map $\omega$ extends to an anti-involution
@@
for $1\leq a,b\leq d$, and a superalgebra homomorphism $\overline{\phi}:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$, given by
${\begin{array}[]{cc}E^{r}_{r^{\prime}}\mapsto-\sum_{a=1}^{d}\partial_{y^{a}_{r}}y_{r^{\prime}}^{a},&E^{r}_{p+s}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\zeta_{s}^{a},\\ E^{p+s}_{r}\mapsto-\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}y_{r}^{a},&E^{p+s}_{p+s^{\prime}}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\zeta_{s^{\prime}}^{a},\\ E^{r}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{x^{a}_{i}},&E^{r}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{\eta^{a}_{j}},\\ E^{p+s}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{x^{a}_{i}},&E^{p+s}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{\eta^{a}_{j}},\\ E^{p+q+i}_{r}\mapsto-\sum_{a=1}^{d}x^{a}_{i}y_{r}^{a},&E^{p+q+i}_{p+s}\mapsto\sum_{a=1}^{d}x^{a}_{i}\zeta_{s}^{a},\\ E^{p+q+m+j}_{r}\mapsto-\sum_{a=1}^{d}\eta^{a}_{j}y_{r}^{a},&E^{p+q+m+j}_{p+s}\mapsto\sum_{a=1}^{d}\eta^{a}_{j}\zeta_{s}^{a},\\ E^{p+q+i}_{p+q+i^{\prime}}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{x^{a}_{i^{\prime}}},&E^{p+q+i}_{p+q+m+j}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{\eta^{a}_{j}},\\ E^{p+q+m+j}_{p+q+i}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{x^{a}_{i}},&E^{p+q+m+j}_{p+q+m+j^{\prime}}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{\eta^{a}_{j^{\prime}}}.\end{array}}$ (4.3)
where $1\leq i,i^{\prime}\leq m$, $1\leq j,j^{\prime}\leq n$, $1\leq r,r^{\prime}\leq p$, and $1\leq s,s^{\prime}\leq q$.
Let $\mathcal{F}$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$ and $\zeta^{a}_{s}$, for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, and $1\leq s\leq q$. With an appropriate inner product, $\mathcal{F}$ may be regarded as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators. The Weyl superalgebra $\mathcal{D}$ acts naturally on $\mathcal{F}$, and so it is a subalgebra of ${\rm End}(\mathcal{F})$.
We simplify notations for the rest of this section. Let
@@
$e_{ab}[k]\mapsto\sum_{i=1}^{m+n}z_{p+q+i}^{k}x^{a}_{i}\partial_{x^{b}_{i}}-\sum_{r=1}^{p+q}z_{r}^{k}y^{b}_{r}\partial_{y^{a}_{r}},$ (4.4)
for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$, and $\overline{\phi}$ induces a superalgebra homomorphism $\overline{\phi}_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$ defined by
@@
for $1\leq r,s\leq p+q$ and $1\leq i,j\leq m+n$.
The Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and ${\mathcal{B}_{p+m|q+n}^{\bf z}}$ are respectively commutative subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$. Thus, the maps $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ and an action of ${\mathcal{B}_{p+m|q+n}^{\bf z}}$, respectively, on the polynomial superalgebra $\mathcal{F}$.
The Cartan anti-involution $\boldsymbol{\theta}$ on ${\mathfrak{gl}}_{d}$, defined in (3.7), gives rise to an anti-involution $\boldsymbol{\theta}:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$, defined by
$\hskip 42.67912pt\boldsymbol{\theta}(A[r])=\boldsymbol{\theta}(A)[r]\qquad\mbox{for $A\in{\mathfrak{gl}}_{d}$ and $r\in{\mathbb{Z}}_{+}$}.$
Proposition 4.1 ([MTV1, Proposition 8.4]).
The anti-involution $\boldsymbol{\theta}:U({\mathfrak{gl}}_{d}[t])\longrightarrow U({\mathfrak{gl}}_{d}[t])$ fixes each element of $\mathcal{B}_{d}^{\bf w}$.
Recall the anti-involution $\omega$ on $\mathcal{D}$ defined in (4.1). It is straightforward to verify that the diagram
[…]
commutes.
Proposition 4.2.
For any ${\boldsymbol{b}}\in\mathcal{B}_{d}^{\bf w}$, $\omega\big{(}\phi_{\bf z}({\boldsymbol{b}})\big{)}=\phi_{\bf z}({\boldsymbol{b}})$.
Proof.
This follows from (4.6) and Proposition 4.1. ∎
4.2. From Berezinians to column determinants
The map $\phi_{\bf z}$ and $\overline{\phi}_{\bf w}$, defined in (4.4) and (4.5), extend naturally to the superalgebra homomorphisms

### v2

@@ 4.1. Column determinants and Berezinians
for some $b_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$.
Let $\widehat{{\mathfrak{z}}}_{p+m|q+n}$ be the subalgebra of $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ generated by
$\{{\rm T}^{r}b_{i}\,|\,i\leq p+m-q-n,\,i\in{\mathbb{Z}},\,r\in{\mathbb{Z}}_{+}\},$
where $\rm T$} is defined by (3.1). According to [MR, Corollary 3.3] (see also [ChL2, Proposition 3.3]), $b_{i}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ for all $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. Since ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$ is $\rm T$}-invariant, we see that $\widehat{{\mathfrak{z}}}_{p+m|q+n}$ is a commutative subalgebra of ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})$.
There is a conjecture asserting that ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{p+m|q+n})=\widehat{{\mathfrak{z}}}_{p+m|q+n}$ (see [MR]). The conjecture holds if $q=n=0$ (see [CM, Theorem 3.1] and also Theorem 3.1). It is also true that ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{1|1})=\widehat{{\mathfrak{z}}}_{1|1}$ ([MM]) and ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{2|1})=\widehat{{\mathfrak{z}}}_{2|1}$ ([AN]). However, it is still unknown whether the conjecture is valid in general.
Now let ${\bf z}=(z_{1},\ldots,z_{p+q+m+n})\in{\mathbb{C}}^{p+q+m+n}$. Define $\mu_{\bf z}\in{\mathfrak{gl}}_{p+m|q+n}^{*}$ by
@@
for $a_{i}\in U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])$ and $s\in{\mathbb{Z}}$. Consider the matrix
${\mathcal{L}}_{{p+m|q+n}}^{\bf z}:=\Big[\Psi^{\bf z}\big(\delta_{i,j}\tau+(-1)^{|i|}E^{i}_{j}[-1]\big)\Big]_{i,j\in\mathbb{I}}=\Big[\delta_{i,j}(\partial_{z}-z_{i})-(-1)^{|i|}E^{i}_{j}(z)\Big]_{i,j\in\mathbb{I}}.$ (4.7)
It is an amply invertible Manin matrix of type $\bf s$} over $U({\mathfrak{gl}}_{p+m|q+n}[t])(\!(z^{-1},\partial_{z}^{-1})\!)$. The following is an immediate consequence of Proposition 4.3.
@@
${\rm Ber}^{{\bf s}^{\sigma}}\big(\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)^{\sigma}\big)={\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big).$
@@
${\rm Ber}^{\bf s}\big({\mathcal{L}}_{{p+m|q+n}}^{\bf z}\big)=\sum_{i=-\infty}^{p+m-q-n}b_{i}^{{\bf z}}(z)\partial_{z}^{i},$ (4.8)
where $b_{i}^{{\bf z}}(z):=\Psi^{\bf z}(b_{i})$. Let ${\mathcal{B}_{p+m|q+n}^{\bf z}}:={\mathcal{B}_{{\mathfrak{gl}}_{p+m|q+n}}^{\bf z}}$ be the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of the series $b_{i}^{{\bf z}}(z)$, for $i\in{\mathbb{Z}}$ with $i\leq p+m-q-n$. It is called the Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$ with respect to $\bf z$}. We remark that $\mathcal{B}_{p+m|q+n}^{\bf z}$} coincides with the subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$ generated by the coefficients of $\Psi^{\bf z}(S)$ for $S\in\widehat{{\mathfrak{z}}}_{p+m|q+n}$ (as $\Psi^{\bf z}({\rm T}(S))=\frac{d}{dz}(\Psi^{\bf z}(S))$). If we take $q=n=0$, we recover the Bethe algebra for ${\mathfrak{gl}}_{p+m}$ by virtue of Proposition 4.5.
According to [MR, Corallary 3.7], the algebra $\mathcal{B}_{p+m|q+n}^{\bf z}$} is a commutative subalgebra of $U({\mathfrak{gl}}_{p+m|q+n}[t])$. It also commutes with the Cartan subalgebra ${\mathfrak{h}}_{{p+m|q+n}}$ of ${\mathfrak{gl}}_{p+m|q+n}$ and hence acts on any weight space of a ${\mathfrak{gl}}_{p+m|q+n}$-module.
For any $\mathcal{B}_{p+m|q+n}^{\bf z}$}-module $V$, let $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{V}$ denote the image of the Bethe algebra $\mathcal{B}_{p+m|q+n}^{\bf z}$} in ${\rm End}(V)$. We call $\big({\mathcal{B}_{p+m|q+n}^{\bf z}}\big)_{V}$ the Bethe algebra for $V$ with respect to $\bf z$}.
4.3. The Bethe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$
Let $\mathcal{F}:={\mathbb{C}}[{\bf x},{\bf y},{\bm{\eta}},{\bm{\zeta}}]$ be the polynomial superalgebra generated by $x^{a}_{i}$, $\eta^{a}_{j}$, $y^{a}_{r}$, and $\zeta^{a}_{s}$, for $i=1,\ldots,m$, $j=1,\ldots,n$, $r=1,\ldots,p$, $s=1,\ldots,q$, and $a=1,\ldots,d$. Here $x^{a}_{i}$ and $y^{a}_{r}$ are even variables while $\eta^{a}_{j}$ and $\zeta^{a}_{s}$ are odd. With an appropriate inner product, $\mathcal{F}$ may be regarded as the Fock space of $d(p+m)$ bosonic and $d(q+n)$ fermionic oscillators (cf. [CL, LZ]). Let $\mathcal{D}:={\mathbb{D}}[{\bf x},{\bf y},{\bm{\eta}},{\bm{\zeta}}]$ be the corresponding Weyl superalgebra, i.e., $\mathcal{D}$ is the associative unital superalgebra generated by $x^{a}_{i},\,\eta^{a}_{j},\,y^{a}_{r},\,\zeta^{a}_{s}$ as well as their derivatives
@@
for $1\leq i\leq m$, $1\leq j\leq n$, $1\leq r\leq p$, $1\leq s\leq q$, and $1\leq a\leq d$. The superalgebra $\mathcal{D}$ acts naturally on $\mathcal{F}$, and so it is a subalgebra of ${\rm End}(\mathcal{F})$.
@@
for $1\leq a,b\leq d$, and a superalgebra homomorphism $\varphi:U({\mathfrak{gl}}_{p+m|q+n})\longrightarrow\mathcal{D}$, given by
$\begin{array}[]{cc}E^{r}_{r^{\prime}}\mapsto-\sum_{a=1}^{d}\partial_{y^{a}_{r}}y_{r^{\prime}}^{a},\quad&E^{r}_{p+s}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\zeta_{s}^{a},\\ E^{p+s}_{r}\mapsto-\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}y_{r}^{a},\quad&E^{p+s}_{p+s^{\prime}}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\zeta_{s^{\prime}}^{a},\\ E^{r}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{x^{a}_{i}},\quad&E^{r}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{y^{a}_{r}}\partial_{\eta^{a}_{j}},\\ E^{p+s}_{p+q+i}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{x^{a}_{i}},\quad&E^{p+s}_{p+q+m+j}\mapsto\sum_{a=1}^{d}\partial_{\zeta^{a}_{s}}\partial_{\eta^{a}_{j}},\\ E^{p+q+i}_{r}\mapsto-\sum_{a=1}^{d}x^{a}_{i}y_{r}^{a},\quad&E^{p+q+i}_{p+s}\mapsto\sum_{a=1}^{d}x^{a}_{i}\zeta_{s}^{a},\\ E^{p+q+m+j}_{r}\mapsto-\sum_{a=1}^{d}\eta^{a}_{j}y_{r}^{a},\quad&E^{p+q+m+j}_{p+s}\mapsto\sum_{a=1}^{d}\eta^{a}_{j}\zeta_{s}^{a},\quad\\ E^{p+q+i}_{p+q+i^{\prime}}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{x^{a}_{i^{\prime}}},\quad&E^{p+q+i}_{p+q+m+j}\mapsto\sum_{a=1}^{d}x_{i}^{a}\partial_{\eta^{a}_{j}},\\ E^{p+q+m+j}_{p+q+i}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{x^{a}_{i}},\quad&E^{p+q+m+j}_{p+q+m+j^{\prime}}\mapsto\sum_{a=1}^{d}\eta_{j}^{a}\partial_{\eta^{a}_{j^{\prime}}}.\end{array}$} (4.10)
where $1\leq r,r^{\prime}\leq p$, $1\leq s,s^{\prime}\leq q$, $1\leq i,i^{\prime}\leq m$, and $1\leq j,j^{\prime}\leq n$. These maps induce a ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-action on $\mathcal{F}$.
Let ${\bm{\pi}}:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ be the *-structure on ${\mathfrak{gl}}_{d}$ defined by
${\bm{\pi}}(e_{ab})=e_{ba}\qquad\mbox{for $1\leq a,b\leq d$.}$
Note that any unitarizable highest weight ${\mathfrak{gl}}_{d}$-module with respect to $\bm{\pi}$} is finite-dimensional. Recall the *-structure ${\bm{\sigma}}_{p}$ on ${\mathfrak{gl}}_{p+m|q+n}$ defined by (3.4). The maps $\bm{\pi}$} and ${\bm{\sigma}}_{p}$ give rise to a *-structure $({\bm{\pi}},{\bm{\sigma}}_{p})$ on ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$. It is well-known that $\mathcal{F}$ decomposes into a direct sum of unitarizable ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{p+m|q+n}$-modules (with respect to $({\bm{\pi}},{\bm{\sigma}}_{p})$). This is called the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{p+m|q+n})$ ([CLZ]). Such unitarizable modules are, in general, infinite-dimensional. Taking $p=q=0$, the maps $\phi$ and $\varphi$ specialize to the superalgebra homomorphisms $U({\mathfrak{gl}}_{d})\longrightarrow{\mathbb{D}}[{\bf x},{\bm{\eta}}]$ and $U({\mathfrak{gl}}_{m|n})\longrightarrow{\mathbb{D}}[{\bf x},{\bm{\eta}}]$, respectively, and the duality reduces to the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m|n})$, which yields a decomposition of ${\mathbb{C}}[{\bf x},{\bm{\eta}}]$ into a direct sum of finite-dimensional ${\mathfrak{gl}}_{d}\times{\mathfrak{gl}}_{m|n}$-modules ([CW, Se1, Se2]). Assuming further that $n=0$, it gives the Howe duality of $({\mathfrak{gl}}_{d},{\mathfrak{gl}}_{m})$ ([H]).
We simplify notations for the rest of this subsection. Let
@@
$e_{ab}[k]\mapsto-\sum_{r=1}^{p+q}z_{r}^{k}y^{b}_{r}\partial_{y^{a}_{r}}+\sum_{i=1}^{m+n}z_{p+q+i}^{k}x^{a}_{i}\partial_{x^{b}_{i}},$ (4.11)
for $1\leq a,b\leq d$ and $k\in{\mathbb{Z}}_{+}$, and $\varphi$ induces a superalgebra homomorphism $\varphi_{\bf w}:U({\mathfrak{gl}}_{p+m|q+n}[t])\longrightarrow\mathcal{D}$ defined by
@@
for $1\leq r,s\leq p+q$, $1\leq i,j\leq m+n$, and $k\in{\mathbb{Z}}_{+}$. Since the Bethe algebras $\mathcal{B}_{d}^{\bf w}$ and $\mathcal{B}_{p+m|q+n}^{\bf z}$} are subalgebras of $U({\mathfrak{gl}}_{d}[t])$ and $U({\mathfrak{gl}}_{p+m|q+n}[t])$, respectively, the maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ induce an action of $\mathcal{B}_{d}^{\bf w}$ on $\mathcal{F}$ and an action of $\mathcal{B}_{p+m|q+n}^{\bf z}$} on $\mathcal{F}$, respectively.
In view of the Howe duality, it is natural to ask if any duality exists in the context of Bethe algebras as well. When $p=q=n=0$, a duality phenomenon is noticed in [TL] (see also [TV]). In [MTV3], the precise statement is made and proved. So far, various generalizations have emerged (see [HM, TU, VY]). We will show that the desired duality exists between $\mathcal{B}_{d}^{\bf w}$ and $\mathcal{B}_{p+m|q+n}^{\bf z}$} in a moment.
The maps $\phi_{\bf z}$ and $\varphi_{\bf w}$ extend naturally to the superalgebra homomorphisms